OpenStudy (ttop0816):

Please help~! Given a test that is normally distributed with a mean of 100 and a standard deviation of 12, find the probability that a sample of 25 scores will have a mean greater than 105. Show your work so I can understand please! :)

OpenStudy (ttop0816):

@agent0smith mind if you help?

OpenStudy (agent0smith):

Ugh sometimes I hate stats. It's easy enough for one test, you just find the z-score and then the probability. But a sample of 25... hmm

OpenStudy (agent0smith):

I *think* you can do it by using the fact 25 tests, with a mean of 100, should have a total score of 25*100 = 2500. The std deviation for the total score of 25 tests would be sqrt25*12 = 60. Those come from the properties of st. dev and mean when multiplying them by a constant. I'm tired and don't really feel like explaining them. And your 25 tests with mean 105 would have a total score 25*105 = 2625 You need P(X > 2625) Now find the z-score for that \[z = \frac{ 2625 - 2500 }{ 60 } = 2.08 \] Now finally, use a z-table to find P(z > 2.08)